OpenStudy (anonymous):

If f(x) varies directly with x and f(x) = 3 when x = 12, then what is the value of f(x) when x = 40? Could someone explain this to me? I got 4 as my answer.

OpenStudy (anonymous):

Didn't I just do this for you haha? Why do you say 4?

OpenStudy (anonymous):

I didn't see that you replied to it! Oops.

OpenStudy (anonymous):

The equation we use for such questions is f(x) = kx where k is a constant, you can use any letter it's a constant. So you just need to find this constant. So what I did earlier, \[f(x) = 3,~~~3=12k \implies k = 3/12 \implies 1/4 \implies f(x) = \frac{ 1 }{ 4 }x\]

OpenStudy (anonymous):

The only possible answers that are listed are 480 160 10 4

OpenStudy (anonymous):

Oh sorry I missed the part that says when x = 40

OpenStudy (anonymous):

So what you have to do here is, we have 1/4, so we have f(x) = 1/4x so now if we have x = 40 we just plug in 40 where the x is, so we get f(x) 40/4 which is?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Do you see what I did?

OpenStudy (anonymous):

Yes! It took me a minute, but it helped that you broke it down for me. Thank you so much!

OpenStudy (anonymous):

Np ^.^